OpenStudy (anonymous):

abcd is a rectangle. if ac = 5x +2 and bd = x + 22, find the value of x

OpenStudy (anonymous):

 The diagonals of a rectangle are equal to each other, so \[AC=BD\] or \[5x+2=x+22\] Now, all you have to do is solve for x. \[5x+2=x+22\]\[4x=20\]\[x=5\]

jhonyy9 (jhonyy9):

do you understand it ?